7. Find the angle between the vectors. ü = (-3,2) and v = (-1,1) e perpendicular. 112 - at
4. Find the vector v such that. ||3|| = 9 and having the same direction as ū. ū=(3,-2)
Answer #1

Giveni u = C-3, &> J = -1, 12 Angle o between rectors iss- Cost = (a) (b) (al (6) u v = 2 (-3(-1) + (2) (1) ( 3+e) . (5) 5(372) To I having find the vector w such same direction that 115 11 = 9 and as I = <3, -27. since I having the same direction as